A time-sharing condominium firm offers prizes to people who visit its projects and listen to a marketing presentation. One prize is a $1,000 savings account. Unfortunately, the account would not be available for forty-five years and requires that the winner pay an initial service fee of $55. If one puts $55 in an investment account, what annual compound rate of return would cause that sum to reach $1,000 in forty-five years?
Answer #1

Use the formula:-

F=P*(1+r)^n

F = 1000

P =55

n= 45

1000= 55(1+r)^45

1000/55 = (1+r)^45

1+r = 1.066

r = 1.066-1

r = 0.066

r = 6.66%